OpenStudy (anonymous):

7x-21/3x+9 divided by 6x-18/4x+12

OpenStudy (anonymous):

\[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a*d}{b*c}\]

OpenStudy (anonymous):